OpenStudy (anonymous):

Write the equation of the line perpendicular to the line x - 5y = -10 and passing through the point (2,5). I really honestly just can't even..

OpenStudy (anonymous):

So one way of writing normal lines \(y=mx+b\) where \(m\) is the slope and \(b\) is the y-intercept. So we know that the first line is actually: \[\eqalign{ &x-5y=-10 \\ &x=5y-10 \\ &x+10=5y \\ &\frac{x+10}{5}=y \\ &0.2x+2=y\\ &y=0.2x+2 }\]

OpenStudy (anonymous):

Ok, how do I figure out the answer?

OpenStudy (anonymous):

So you know that the line has a slope of \(0.2\) If the slope of the original line was \(m\), the slope of the perpendicular line is \(-\frac{1}{m}\). This means that the slope of the perpendicular line is: \(-\frac{1}{0.2}=-5\). So so far the equation of the perpendicular line is: \[y=-5x+b\] We need to find \(b\). Lets sub in a point such as the one we are given \((2,5)\): \[\eqalign{ &5=-5(2)+b \\ &5=-10+b \\ &5+10=b \\ &15=b \\ }\] So therefore the equation is: \[y=-5x+15\] Or a version which looks nicer: \[y=15-5x\]

OpenStudy (anonymous):

Can you help me with a couple more?

OpenStudy (anonymous):

Ehh...Maybe...Yeah I could. Where?